OpenStudy (anonymous):

-5a(2b+3c) simplify please HELLLPPPPPPPPPPPPPPPPPPP

OpenStudy (luigi0210):

Welcome to Openstudy, first what you have to do is distribute that -5a.

OpenStudy (luigi0210):

We are not here to give you answers. you have to do the work.

OpenStudy (anonymous):

how to solve parenthis 1.when there is positive sign before a bracket then no change of sign in the bracket. for example +(2x-3y+4)=+2x-3y+4 2. when there is negative sign before a bracket then sign inside the bracket changes. for example -(2x-3y+4)=-2x+3y-4 3. when there is a positive number before a bracket then number is multiplied and sign remains the same. for example 2t (2x-3y+4)=4tx-6ty+8t note 2t(2x-3y+4)is same as +2t(2x-3y+4) 4when there is a negative number before a bracket then number is multiplied and sign inside the bracket changes. for example -2t(2x-3y+4)=-4tx+6ty-8t

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."